Cane Company manufactures two products called Alpha and Beta that sell for $120 and $80, respectively. Each product uses only one type of raw material that costs $6 per pound. The company has the capacity to annually produce 100,000 units of each product. Its average cost per unit for each product at this level of activity are given below:

Alpha Beta
Direct materials $ 30 $ 12
Direct labor 20 15
Variable manufacturing overhead 7 5
Traceable fixed manufacturing overhead 16 18
Variable selling expenses 12 8
Common fixed expenses 15 10
Total cost per unit $ 100 $ 68

The company considers its traceable fixed manufacturing overhead to be avoidable, whereas its common fixed expenses are unavoidable and have been allocated to products based on sales dollars.

1.1. What is the total amount of traceable fixed manufacturing overhead for each of the two products?

Alpha Beta
Traceable fixed manufacturing overhead

2. What is the company’s total amount of common fixed expenses?

3. Assume that Cane expects to produce and sell 80,000 Alphas during the current year. One of Cane's sales representatives has found a new customer who is willing to buy 10,000 additional Alphas for a price of $80 per unit. What is the financial advantage (disadvantage) of accepting the new customer's order?
Answer #1
1.1
Alpha Beta
Traceable fixed manufacturing overhead per unit 16 18
X Units produced 100000 100000
Traceable fixed manufacturing overhead 1600000 1800000
2
Total amount of common fixed expenses 2500000 =100000*(15+10)
3
Direct materials 30
Direct labor 20
Variable manufacturing overhead 7
Variable selling expenses 12
Unit variable cost 69
Incremental revenue from sales 800000 =10000*80
Less : Incremental costs 690000 =10000*69
Net change in income 110000
Financial advantage 110000